Au – 198 is used in the diagnosis of liver problem. The half- life of Au -198 is 2.69 days.


If you begin with 2.8 µg of this gold isotope, what mass remains in 10.8 days?


1
Expert's answer
2022-02-23T06:04:04-0500

Solution:

The half-life (t1/2) of Au-198 is 2.69 days.


From law of radioactive decay:

t1/2 = ln(2) / λ

where λ = decay constant

Therefore,

λ = ln(2) / t1/2

λ = (0.693) / (2.69 days) = 0.2576 days−1


Radioactive decay follows the kinetics of a first order reaction:

ln[Au-198]t = −λt + ln[Au-198]o


[Au-198]t = unknown

[Au-198]o = 2.8 µg

λ = 0.2576 days−1

t = 10.8 days


Therefore,

ln[Au-198]t = −(0.2576 days−1 × 10.8 days) + ln(2.8)

ln[Au-198]t = −2.78208 + 1.02962

ln[Au-198]t = −1.75246

[Au-198]t = e−1.75246 = 0.1733 µg = 0.17 µg

[Au-198]t = 0.17 µg


Answer: 0.17 µg of Au-198 remains